OpenStudy (anonymous):

every morning jim jogs from his house to a nearby park and back. on thursday jim jogged 8 mph on the way to the park and 4 mph on the way back home. jim's total jog took him 1.5 hours. how long did it take jim to jog from his house to the park? how long did it take him to jog from the park home?

OpenStudy (anonymous):

Try drawing a picture first, then set up some distance = rate × time equations.

OpenStudy (anonymous):

does the question give you the distance he traveled?

OpenStudy (anonymous):

the speed of jogging is twice when jim goes to park than that of returning, then the time taken to go to park will be half of the time taken to return total time is 1.5 hours so, time to go to park= .5 hours=30 minutes fyi, this was posted by user Arnab09 11 months ago.

OpenStudy (anonymous):

This one is pretty easy to guess-and-check at using proportional thinking. The set up is to use d=r*t twice. The distance, d, is the same both ways, so r1*t1 = r2*t2 The total time is 1.5 hours, so t1 + t2 = 1.5 r1 and r2 are given, so this can be solved with substitution.